OpenStudy (anonymous):

The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples.

OpenStudy (wolf1728):

If they are Pythagorean Triples, then the highest number squared should equal the sum of the squares of the other 2 numbers. 25^2 = 7^2 + 24^2 = 49 + 576 =625 41^2 = 9^2 + 40^2 = 81 + 1,600 = 1,681 They are both Pythagorean Triples.